OpenStudy (anonymous):

Solve the linear system by using combination(elimination) 0.2x-0.5y=-3.8 0.3x+0.4y=10.4

OpenStudy (anonymous):

(0.2x-0.5y=-3.8) 0.4 (0.3x+0.4y=10.4) 0.5 --------------- .08x - 0.2y = -1.52 0.15x + 0.2y = 5.2 ------------------ .23x = 6.72 (divide each side by .23 to isolate x)] x = 29.21 (rounded it to the nearest hundredths)

OpenStudy (anonymous):

0.2y is eliminated